# Three Cars Approaching within 100m! Enhancing Distant Geometry by Tri-Axis Voxel Scanning for Camera-based Semantic Scene Completion

### v1

@@ 2 Related Work
To address these issues, we propose a model incorporating the Scan Module applying masked self-attention with axis-specific masks and the Scan Loss function, which extends cross-entropy loss.
@@ 3.1 Preliminary: Distance-Dependent Completion Imbalance in Camera-Based SSC
From these analyses, we gain the insight that the prediction accuracy of current camera-based SSC methods tends to decrease as the distance from the viewpoint increases. To tackle these challenges, we introduce ScanSSC, designed to execute near-to-far geometric refinement in an axis-wise manner. For clarity, we define the terms as follows: “near-to-far” refers to front-to-back along the depth axis, center-to-side along the width axis, and high-to-low along the height axis. For the height axis, a BEV serves as the reference for the definition, as the complexity of GT geometry is inversely related to height; lower positions are more challenging to complete due to a greater concentration of object voxels compared to higher positions.
@@ 3.3 Scan Module
Scan Block. Like modern vision Transformer blocks [21, 36, 37], the Scan block is composed of a sequence of self-attention (SA) and feed-forward network (FFN) subblocks, each featuring pre-normalization and a residual connection. For the SA layer, the Scan block utilizes masked self-attention [26] with an axis-specific mask $\mathbf{M}_{axis}\in\{\mathbf{M}_{dep},\mathbf{M}_{wid},\mathbf{M}_{hgt}\}$, where each mask corresponds to the depth, width, and height axes, respectively. We specifically organize the masking positions of $\mathbf{M}_{dep},\mathbf{M}_{wid}$, and $\mathbf{M}_{hgt}$ based on the analysis in Sec. 3.1.
$\mathbf{M}_{dep},\mathbf{M}_{wid}$, and $\mathbf{M}_{hgt}$ are designed to enable distant voxel features to reference preceding voxel features while simultaneously preventing the reverse, thereby minimizing the influence of inaccurate features from distant voxels on previous voxels for each axis. Consequently, we apply a cascading mask that enables attention computation for preceding voxels while blocking subsequent voxels. Since the indices for the depth, width, and height axes of the voxel grid are arranged from front to back, left to right, and bottom to top, $\mathbf{M}_{dep}$ is configured as an upper triangular matrix with zeros along the diagonal. In contrast, $\mathbf{M}_{wid}$ is designed in an hourglass shape, while $\mathbf{M}_{hgt}$ resembles a lower triangular matrix with a zero diagonal. Additionally, we define a margin region within a specific range of well-established near-viewpoint voxels, where masking is removed to allow unrestricted interactions. For $\mathbf{M}_{dep}$, $\mathbf{M}_{wid}$, and $\mathbf{M}_{hgt}$, the margin regions are by default set from the start of “near-to-far” of each axis as follows: 50% backward along the depth axis, 25% to each side (for a total of 50%) along the width axis, and 0% along the height axis.
The 3D feature $\mathbf{F}^{3D}$, obtained through viewing transformation, is separated into three different flattened features: $\mathbf{F}^{3D}_{dep}\in\mathbb{R}^{(\hat{Y}\hat{Z})\times\hat{X}\times C}$, $\mathbf{F}^{3D}_{wid}\in\mathbb{R}^{(\hat{X}\hat{Z})\times\hat{Y}\times C}$, and $\mathbf{F}^{3D}_{hgt}\in\mathbb{R}^{(\hat{X}\hat{Y})\times\hat{Z}\times C}$. Each of the three features is individually input into a corresponding Scan block along with the respective attention masks $\mathbf{M}_{dep},\mathbf{M}_{wid}$, and $\mathbf{M}_{hgt}$. Each per-axis Scan block operates as follows:
@@
To calculate the voxel-wise weight values, $\mathbf{W}\in\mathbb{R}^{\hat{X}\times\hat{Y}\times\hat{Z}\times 3}$, these features are concatenated along the channel dimension, followed by a linear projection and then a softmax function, as shown below:
$\mathbf{W}=\text{Softmax}(\text{Linear}(\text{Concat}(\mathbf{\hat{F}}^{3D}_{dep},\mathbf{\hat{F}}^{3D}_{wid},\mathbf{\hat{F}}^{3D}_{hgt}))).$ (5)
@@
$\mathbf{\hat{F}}^{3D}=\sum_{i}^{3}\mathbf{W}_{[:,:,:,i]}\otimes\mathbf{\hat{F}}^{3D}_{i},$ (6)
where $\mathbf{\hat{F}}^{3D}_{i}$ is an element of the set $\{\mathbf{\hat{F}}^{3D}_{dep},\mathbf{\hat{F}}^{3D}_{wid},\mathbf{\hat{F}}^{3D}_{hgt}\}$ and the operation $\otimes$ represents element-wise multiplication.

### v2

@@ 2 Related Work
To address these issues, we propose a model incorporating the Scan Module applying masked self-attention with axis-specific masks and the Scan Loss function, which extends cross-entropy loss [47].
@@ 3.1 Preliminary: Distance-Dependent Completion Imbalance in Camera-Based SSC
From these analyses, we gain the insight that the prediction accuracy of current camera-based SSC methods tends to decrease as the distance from the viewpoint increases. To tackle these challenges, we introduce ScanSSC, designed to execute near-to-far geometric refinement in an axis-wise manner. For clarity, we define the terms as follows: “near-to-far" refers to front-to-back along the depth axis, center-to-side along the width axis, and high-to-low along the height axis. For the height axis, a BEV serves as the reference for the definition, as the complexity of GT geometry is inversely related to height; lower positions are more challenging to complete due to a greater concentration of object voxels compared to higher positions.
@@ 3.3 Scan Module
Scan Block. Like modern vision Transformer blocks [24, 43, 44], the Scan block is composed of a sequence of self-attention (SA) and feed-forward network (FFN) subblocks, each featuring pre-normalization and a residual connection. For the SA layer, the Scan block utilizes masked self-attention [33] with an axis-specific mask $\textbf{M}_{axis}\in\{\textbf{M}_{dep},{\textbf{M}}_{wid},{\textbf{M}}_{hgt}\}$, where each mask corresponds to the depth, width, and height axes, respectively. We specifically organize the masking positions of $\textbf{M}_{dep},\textbf{M}_{wid}$, and $\textbf{M}_{hgt}$ based on the analysis in Sec. 3.1.
$\textbf{M}_{dep},\textbf{M}_{wid}$, and $\textbf{M}_{hgt}$ are designed to enable distant voxel features to reference preceding voxel features while simultaneously preventing the reverse, thereby minimizing the influence of inaccurate features from distant voxels on previous voxels for each axis. Consequently, we apply a cascading mask that enables attention computation for preceding voxels while blocking subsequent voxels. Since the indices for the depth, width, and height axes of the voxel grid are arranged from front to back, left to right, and bottom to top, $\textbf{M}_{dep}$ is configured as an upper triangular matrix with zeros along the diagonal. In contrast, $\textbf{M}_{wid}$ is designed in an hourglass shape, while $\textbf{M}_{hgt}$ resembles a lower triangular matrix with a zero diagonal. Additionally, we define a margin region within a specific range of well-established near-viewpoint voxels, where masking is removed to allow unrestricted interactions. For $\textbf{M}_{dep}$, $\textbf{M}_{wid}$, and $\textbf{M}_{hgt}$, the margin regions are by default set from the start of “near-to-far” of each axis as follows: 50% backward along the depth axis, 25% to each side (for a total of 50%) along the width axis, and 0% along the height axis.
The 3D feature $\textbf{F}^{3D}$, obtained through viewing transformation, is separated into three different flattened features: $\textbf{F}^{3D}_{dep}\in\mathbb{R}^{(\hat{Y}\hat{Z})\times\hat{X}\times C}$, $\textbf{F}^{3D}_{wid}\in\mathbb{R}^{(\hat{X}\hat{Z})\times\hat{Y}\times C}$, and $\textbf{F}^{3D}_{hgt}\in\mathbb{R}^{(\hat{X}\hat{Y})\times\hat{Z}\times C}$. Each of the three features is individually input into a corresponding Scan block along with the respective attention masks $\textbf{M}_{dep},\textbf{M}_{wid}$, and $\textbf{M}_{hgt}$. Each per-axis Scan block operates as follows:
@@
To calculate the voxel-wise weight values, $\textbf{W}\in\mathbb{R}^{\hat{X}\times\hat{Y}\times\hat{Z}\times 3}$, these features are concatenated along the channel dimension, followed by a linear projection and then a softmax function, as shown below:
$\textbf{W}=\text{Softmax}(\text{Linear}(\text{Concat}(\mathbf{\hat{F}}^{3D}_{dep},\mathbf{\hat{F}}^{3D}_{wid},\mathbf{\hat{F}}^{3D}_{hgt}))).$ (5)
@@
$\mathbf{\hat{F}}^{3D}=\sum_{axis}^{\{dep,wid,hgt\}}\mathbf{W}_{[:,:,:,axis]}\otimes\mathbf{\hat{F}}^{3D}_{axis},$ (6)
where $\otimes$ represents element-wise multiplication.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2023R1A2C200337911 and No. RS-2023-00220762).
[…]
Supplementary Material
A Dataset and Metric
Dataset. We evaluate ScanSSC on SemanticKITTI [3] and SSCBench-KITTI-360 [17] datasets. SemanticKITTI is derived from the KITTI Odometry [7] benchmark, consisting of 22 outdoor scenes captured by LiDAR scans and stereo images. These 22 scenes are split into 10 training scenes, 1 validation scene, and 11 test scenes. The ground truth voxel grids have dimensions of 256$\times$256$\times$32, with each voxel measuring (0.2m, 0.2m, 0.2m), annotated with 21 semantic classes (19 semantics, 1 empty and 1 unknown). SSCBench-KITTI-360 is extracted from the KITTI-360 [19], comprising 7 training scenes, 1 validation scene, and 1 test scene. It includes 19 semantic classes (18 semantics and 1 free).
Metric. Following standard practices in related works [16, 12, 45, 13], we use the mean Intersection over Union (mIoU) to assess the overall performance of semantic scene completion (SSC) and Intersection over Union (IoU) to evaluate the performance of semantic-agnostic scene completion.
B More Details
Implementation Details. We train ScanSSC for 25 epochs on 4 NVIDIA A6000 GPUs with a batch size of 4. The AdamW optimizer [25] is used with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a maximum learning rate of $3\times 10^{-4}$. For the learning rate schedule, we employ a multi-step scheduler, reducing the learning rate by a factor of 0.1 at the 20th epoch.
Architectural Details. Similar to related works [4, 11, 38, 45], we employ a 2D UNet image encoder built upon a pretrained EfficientNetB7 [32]. Following previous stereo-based methods [12, 16, 45], we utilize the MobileStereoNet [30] as the depth estimator. In the viewing transformation, we adopt the depth network from CGFormer [45], which modifies the BEVDepth [15]. There are 3 deformable attention layers for cross-attention and 2 for self-attention, with 8 sampling points per reference point in both heads. The spatial mixing network consists of 3 stages, each with 2 residual blocks [8].
C Computational Cost
We report the computational cost of ScanSSC compared to CGFormer [45] in Tab. C.1. ScanSSC shows competitive efficiency, with only a slight increase in parameters and inference time. However, it achieves notable performance improvements, with a 0.13 increase in IoU and a 0.77 increase in mIoU on the SemanticKITTI test set, highlighting the effectiveness of our method.
Method Params (M) Inference Time (ms) IoU mIoU
CGFormer 122 566 44.41 16.63
ScanSSC 145 674 44.54 17.40
Table C.1: Comparison of computational costs with CGFormer [45]. The inference time for a single sample of SemanticKITTI [3] validation set is measured on 1 NVIDIA A6000 GPU.
D Additional Ablation Studies
We provide additional ablation studies to evaluate the effectiveness of the subcomponents of ScanSSC. Consistent with the manuscript, all experiments are conducted on the SemanticKITTI [3] validation set.
[…]
Method IoU mIoU
Concat $\rightarrow$ Linear 45.90 16.32
Average 46.17 16.51
Weighted Sum 45.90 16.28
Tri-Feature Fusion 45.95 17.12
Table D.1: Ablation study on the tri-feature fusion method of ScanSSC.
Loss Scaling Coefficient for Scan Loss. We conduct an ablation study on the loss scaling coefficient of the Scan Loss, $\mathcal{L}_{scan}$, as shown in Fig. D.1. When the coefficient is set to 1, the highest mIoU score of 17.12 is observed, while it decreases as the coefficient moves further from 1 overall. We find that the training mIoU consistently increases proportionally with $\lambda_{scan}$ throughout the entire training. From this result, we infer that an excessively high value of $\lambda_{scan}$ can lead to overfitting of the model, highlighting the importance of selecting an appropriate scaling coefficient.
Figure D.1: Performance comparison by loss scaling coefficient for Scan Loss.
Scan Loss vs. General Cross-Entropy. Since the proposed Scan Loss is equivalent to the cross-entropy [47] of the cumulatively averaged logit, incorporating it can be seen as analogous to either modifying the distribution of the loss coefficient for the existing voxel-wise cross-entropy or simply increasing its scale. Hence, we compare the performance of ScanSSC when it is substituted with a simple coefficient adjustment for the voxel-wise cross-entropy loss, as shown in Tab. D.2. For (a), to assign higher weights to distant voxels, we first generate a bilinearly interpolated weight along each axis, ranging from 0 to 1 in the corresponding near-to-far direction, then use the average of these weights as the coefficient for the existing cross-entropy loss ($\lambda_{tri}$). For (b), we simply apply a higher scalar weight ($\lambda_{ce}$) to the voxel-wise cross-entropy loss. Here, we set $\lambda_{ce}$ to 10, as the converged loss scales of both methods are similar.
Method IoU mIoU
(a) $\lambda_{tri}$$\mathcal{L}_{ce}$ 46.15 16.50
(b) $\lambda_{ce}$$\mathcal{L}_{ce}$ 45.04 16.74
$\mathcal{L}_{ce}$ + $\mathcal{L}_{scan}$ (Ours) 45.95 17.12
Table D.2: Performance comparison between incorporating Scan Loss and adjusting the coefficient for voxel-wise cross-entropy loss [47]. The other training losses remain unchanged during training.
Comparing with (a), we observe that Scan Loss significantly enhances mIoU by leveraging the semantic distribution of previous voxels to transmit signals to the target voxel. This demonstrates that instead of merely assigning higher weights to distant voxels, utilizing the semantic distribution of previous voxels to propagate signals more effectively is a superior approach. In addition, when compared to (b), the result demonstrates that the simple increase in the weight of the existing cross-entropy does not lead to performance improvement, as it results in significantly lower IoU and mIoU scores.
[…]
Method Spatial-mixing net. Tri-feature fusion IoU mIoU
(a) 44.91 15.90
(b) 45.90 16.32
(c) 45.08 16.11
ScanSSC 45.95 17.12
Table D.3: Ablation study on the subsidiary components of the ScanSSC.
Ablation Study of Using Tri-Axes Features.
Figure D.2: Visualization and comparison of the results of applying the Scan Module and Scan Loss to each individual axis and ScanSSC on the SemanticKITTI [3] validation set.
To demonstrate the validity of using features from all three axes, we visualize and compare the results obtained using features from each axis with those obtained through ScanSSC. The results obtained using each axis individually are represented as ‘Depth Only,’ ‘Width Only,’ and ‘Height Only,’ and are shown in Fig. D.2. Overall, ScanSSC, which utilizes features from all three axes, demonstrates significantly more plausible results. Using features from only a single axis tends to result in inaccurate predictions for distant vehicles, side road areas, and occluded objects. In contrast, ScanSSC, which combines features from all three axes, achieves significantly more reliable and accurate predictions. We hypothesize that this improvement arises from the complementary nature of features from each axis, which together enable a more comprehensive understanding of the entire scene.
E Analysis
Quantitative Result for Distant Geometry. This study aims to improve the overall reconstruction in distant regions. To support this, quantitative analysis results are presented in Tab.2 of the manuscript. Additionally, for a more detailed analysis, Tab. E.1 provides group-wise mIoU for distant $1/2$ regions along each axis ($1/4$ on both sides for the width axis), following the categorization from the official SemanticKITTI [3] website.
Large class group Small class group
Method Axis Ground Structure Nature Total Vehicle Human Object Total
CGFormer Dep. 24.87 17.66 19.57 21.98 6.92 0.07 2.86 3.95
ScanSSC Dep. 26.29 17.89 19.23 22.60 7.03 0.19 3.19 4.11
CGFormer Wid. 14.23 15.81 16.51 15.28 1.24 0.15 1.35 0.97
ScanSSC Wid. 16.26 14.65 15.96 15.95 1.39 0.57 1.49 1.19
CGFormer Hgt. 29.90 23.15 25.73 27.49 13.23 2.79 6.73 8.61
ScanSSC Hgt. 31.69 21.50 25.91 28.25 13.78 3.44 7.11 9.14
Table E.1: Per-group mIoUs on the SemanticKITTI [3] validation set.
This result highlights ScanSSC’s effectiveness in distant regions, demonstrating its superior performance on both large and small geometries, particularly outperforming CGFormer [45] in all small class groups.
Figure E.1: Visualization results of the non-axis-aligned cases of CGFormer [45] and ScanSSC on the SemanticKITTI [3] validation set.
Figure E.2: Visualization results of CGFormer [45] and ScanSSC under various conditions (e.g., shadow, occlusion) on the SemanticKITTI [3] validation set.
Analysis of the Non-Axis-Aligned Cases. Since the proposed Scan Module and Scan Loss operate axis-wise, ScanSSC is effective in most axis-aligned driving scenarios, as demonstrated by the qualitative results in the manuscript. However, this raises the question of whether ScanSSC’s operation might be less effective in non-axis-aligned scenes. To investigate this, we conduct additional evaluations of ScanSSC in non-axis-aligned scenarios. Since the SemanticKITTI benchmark dataset does not explicitly categorize curve road scenes, we manually classify these cases. Numerically, ScanSSC significantly outperforms CGFormer, achieving a mIoU of 14.56 and an IoU of 42.38, compared to CGFormer’s 13.77 and 41.96. As shown in Fig. E.1, ScanSSC performs comparably overall without side effects. Specifically, its performance is on par for small objects; however, it reconstructs roads significantly better in distant regions.
Analysis of the Various Scene Conditions. We conduct additional analyses to demonstrate the superiority of ScanSSC under various conditions (e.g., shadow, occlusion). Since existing benchmark datasets for SSC do not categorize various environments, we manually filter shady and highly occluded scenarios from the RGB images of the SemanticKITTI dataset.
As shown in Fig. E.2, in the shady scenario, ScanSSC clearly distinguishes both nearby and distant vehicles covered by shadows, whereas CGFormer fails to do so. In addition, in the occluded scenario, ScanSSC effectively reconstructs the right-side road obscured by nearby vehicles and accurately identifies distant cars that are partially occluded. These results demonstrate ScanSSC’s robustness across diverse and challenging scenes.
F Additional Qualitative Results
Figure F.2: More qualitative comparison results on the SemanticKITTI [3] validation set.
We present additional qualitative comparisons with VoxFormer [16] and CGFormer [45], as visualized in Fig. F.2. These results are randomly selected from the SemanticKITTI [3] validation set.
G Pytorch-like Pseudocode of Scan Module and Scan Loss
To facilitate a comprehensive understanding of the proposed Scan Module and Scan Loss, we present the PyTorch-like [28] pseudocode for each in Algorithm G.1 and Algorithm G.2, respectively.
Algorithm G.1 PyTorch Style Pseudocode of Scan Module.
import torch
import torch.nn as nn
class ScanModule(nn.Module):
def __init__(self, dim):
# declare axis-specific Scan Blocks
self.dep_block = ScanBlock(dim)
self.wid_block = ScanBlock(dim)
self.hgt_block = ScanBlock(dim)
def forward(self, x):
X_, Y_, Z_, C = x.size()
x_dep = x.permute(1, 2, 0, 3).flatten(0,1)
x_wid = x.permute(0, 2, 1, 3).flatten(0,1)
x_hgt = x.flatten(0,1)
# axis-specific masks
dep_attn_mask = torch.triu(torch.ones(X_, X_), diagonal=1)==1
dep_attn_mask[:, :X_//2] = False # depth-axis margin region
wid_attn_mask = torch.tril(torch.ones(Y_//2, Y_//2), diagonal=-1)==1
wid_attn_mask = torch.cat((wid_attn_mask, wid_attn_mask.flip(dim=[-1])), dim=-1)
wid_attn_mask = torch.cat((wid_attn_mask, wid_attn_mask.flip(dims=[-2])), dim=0)
wid_attn_mask[:, Y_//4:-(Y_//4)] = False # width-axis margin region
hgt_attn_mask = torch.tril(torch.ones(Z_), diagonal=-1)==1
# axis-wise voxel scanning
x_dep = self.dep_block(x_dep, dep_attn_mask)
x_wid = self.wid_block(x_wid, wid_attn_mask)
x_hgt = self.hgt_block(x_hgt, hgt_attn_mask)
x_dep = x_dep.reshape(Y_, Z_, X_, C).permute(2, 0, 1, 3)
x_wid = x_wid.reshape(X_, Z_, Y_, C).permute(0, 2, 1, 3)
x_hgt = x_hgt.reshape(X_, Y_, Z_, C)
return x_dep, x_wid, x_hgt
class ScanBlock(nn.Module):
def __init__(self, dim):
self.norm1 = nn.LayerNorm(dim)
self.masked_sa = nn.MultiheadAttention(dim)
self.norm2 = nn.LayerNorm(dim)
self.ff1 = nn.Linear(dim, dim*2)
self.activation = nn.ReLU()
self.ff2 = nn.Linear(dim*2, dim)
def forward(self, x, attn_mask):
B_, L_, C = x.size()
# Masked Self-Attention
x_norm1 = self.norm1(x)
x = x + self.masked_sa(x_norm1, x_norm1, x_norm1,
attn_mask = attn_mask)
# Feed Forward Network
x_norm2 = self.norm2(x)
x = x + self.ff2(self.activation(self.ff1(x_norm2)))
return x
Algorithm G.2 PyTorch Style Pseudocode of $\mathcal{L}_{\text{scan}}$.
import torch
import torch.nn.functional as F
def ScanLoss(logit, target):
P, X_, Y_, Z_ = logit.size()
# back to front
cum_x = torch.cumsum(logit.flip((1)), axis=1)
# sides to center
cum_y_l = torch.cumsum(logit[:Y_//2], axis=2)
cum_y_r = torch.cumsum(logit[Y_//2:].flip((2)), axis=2)
cum_y = torch.cat([cum_y_l, cum_y_r], dim=2)
# bottom to top
cum_z = torch.cumsum(logit, axis=3)
# to logit value scaling
cum_x /= torch.arange(1, X_+1)
cum_y /= torch.arange(1, Y_+1)
cum_z /= torch.arange(1, Z_+1)
# same with logits
X_, Y_, Z_ = target.size()
target = F.one_hot(target).permute(3, 0, 1, 2)
cum_x_t = torch.cumsum(target.flip((1)), dim=1)
cum_y_l_t = torch.cumsum(target[:Y_//2], dim=2)
cum_y_r_t = torch.cumsum(target[Y_//2:].flip((2)), dim=2)
cum_y_t = torch.cat([cum_y_l_t, cum_y_r_t], axis=2)
cum_z_t = torch.cumsum(target, dim=3)
cum_x_t /= torch.arange(1, X_+1)
cum_y_t /= torch.arange(1, Y_+1)
cum_z_t /= torch.arange(1, Z_+1)
L_scan_x = F.cross_entropy(cum_x, cum_x_t, reduction=’mean’)
L_scan_y = F.cross_entropy(cum_y, cum_y_t, reduction=’mean’)
L_scan_z = F.cross_entropy(cum_z, cum_z_t, reduction=’mean’)
L_scan = L_scan_x + L_scan_y + L_scan_z
return L_scan